# Understanding the Nature of Depth-1 Equivariant Quantum Circuit

### v1

@@ 1.2 Contributions
  • A size-invariant understanding of parameters: We provide a new explanation for the roles of the trainable parameters ($\gamma$ and $\beta$) at Depth 1 of the EQC, extending beyond the concept of equivariance in [Skolik2022EquivariantQC]. By analyzing the closed-form expressions of the Depth 1 EQC [akshay2021parameter, analytical-qaoa, exp-qaoa-ising, qaoa-ferm], we show how these parameters govern effective QRL policies for TSP. On one hand, while QAOA uses variational optimization to minimize Hamiltonian expectations, QRL (using Q-Learning) involves optimizing the Temporal-Difference Loss [sutton-barto-2018], with Q-values derived from ZZZZ expectations and guided by a Markov Decision Process(MDP). This fundamental methodological difference motivates our analysis of Depth 1 EQCs in reinforcement learning.
  • Size-invariant trainability:. As a consequence of the size-invariant interpretation of parameters, we conclude that effective QRL policies with Depth 1 EQCs reside in a compact, problem size-invariant parameter region. Leveraging this, we propose the Size-Invariant Grid Search (SIGS), which reduces training time for EQC-based QRL by 27 times (e.g., 151 minutes (using the Analytical Expression) to 5.43 minutes for TSP with 100 nodes), and enables efficient simulation of instances with up to 350 nodes (longest run completes under 140140 minutes). This allows us to analyze scalability of Depth 1 EQCs well beyond the tractable limits of classical quantum simulation.
@@ 2.1 Markov Decision Process for TSP
    Episodes and Steps: An episode represents a complete TSP tour. The initial step $t=0$ is the initial state the RL agent is in, denoted by $s_{0}$. Without loss of generality, our tours begin with the node at index 11 (assuming one-indexed), where only node 1 is visited. The terminal step $T$ is the step |V|1|V|-1, which is when all nodes are visited, and $t$ being some permutation of the $|V|$ vertices.
@@ 2.2 Parameterized Quantum Circuit
The Equivariant Quantum Circuit (EQC) is an example of an PQC. A diagram of the EQC is provided in Figure LABEL:fig:eqc-diagram for a $n=4$ qubit system. In the formulation of the EQC, there is a one-to-one mapping between qubits and TSP nodes, which means an $n$ qubit system is only able to handle TSP instances of size $n$. There are two kinds of parameterized gates in the EQC as shown: a ZZZZ-gate and a $R_{x}$ gate. Note that $\gamma$ is shared across all ZZZZ-gates, while $\beta$ is shared across all $R_{x}$ gates. Measurements are recorded with the $ZZ=Z\bigotimes Z$ observable. The analytical expression for the output of the EQC in the basis of the observable $\langle Z_{t}Z_{a}\rangle$, where $t$ is the last node of the present tour and $a$ is an unexplored next node, is given in [Skolik2022EquivariantQC]:

### v2

@@ 1.2 Contributions
  • A size-invariant understanding of parameters: We provide a new explanation for the roles of the trainable parameters ($\gamma$ and $\beta$) at Depth 1 of the EQC, extending beyond the concept of equivariance in [Skolik2022EquivariantQC]. By analyzing the closed-form expressions of the Depth 1 EQC [akshay2021parameter, analytical-qaoa, exp-qaoa-ising, qaoa-ferm], we show how these parameters govern effective QRL policies for TSP. On one hand, while QAOA uses variational optimization to minimize Hamiltonian expectations, QRL (using Q-Learning) involves optimizing the Temporal-Difference Loss [sutton-barto-2018], with Q-values derived from $ZZ$ expectations and guided by a Markov Decision Process(MDP). This fundamental methodological difference motivates our analysis of Depth 1 EQCs in reinforcement learning.
  • Size-invariant trainability:. As a consequence of the size-invariant interpretation of parameters, we conclude that effective QRL policies with Depth 1 EQCs reside in a compact, problem size-invariant parameter region. Leveraging this, we propose the Size-Invariant Grid Search (SIGS), which reduces training time for EQC-based QRL by 27 times (e.g., 151 minutes (using the Analytical Expression) to 5.43 minutes for TSP with 100 nodes), and enables efficient simulation of instances with up to 350 nodes (longest run completes under $140$ minutes). This allows us to analyze scalability of Depth 1 EQCs well beyond the tractable limits of classical quantum simulation.
@@ 2.1 Markov Decision Process for TSP
    Episodes and Steps: An episode represents a complete TSP tour. The initial step $t=0$ is the initial state the RL agent is in, denoted by $s_{0}$. Without loss of generality, our tours begin with the node at index $1$ (assuming one-indexed), where only node 1 is visited. The terminal step $T$ is the step |V|1|V|-1, which is when all nodes are visited, and $t$ being some permutation of the $|V|$ vertices.
@@ 2.2 Parameterized Quantum Circuit
The Equivariant Quantum Circuit (EQC) is an example of an PQC. A diagram of the EQC is provided in Figure LABEL:fig:eqc-diagram for a $n=4$ qubit system. In the formulation of the EQC, there is a one-to-one mapping between qubits and TSP nodes, which means an $n$ qubit system is only able to handle TSP instances of size $n$. There are two kinds of parameterized gates in the EQC as shown: a $ZZ$-gate and a $R_{x}$ gate. Note that $\gamma$ is shared across all $ZZ$-gates, while $\beta$ is shared across all $R_{x}$ gates. Measurements are recorded with the $ZZ=Z\bigotimes Z$ observable. The analytical expression for the output of the EQC in the basis of the observable $\langle Z_{t}Z_{a}\rangle$, where $t$ is the last node of the present tour and $a$ is an unexplored next node, is given in [Skolik2022EquivariantQC]: